Director's Time

Observers: B.Rothberg, O. Kuhn, M. (Tucson)

ISA: O.Kuhn, B.Rothberg (Tucson)
Telescope Operator: G.Bechetti (LBTO)
AO Support: J.Christou
Telescope Support: K.Newton (LBTO)
Software Support: C.Biddick
Instrument Support: E.Shield (LBTO)
Engineering Support: M.Gusick

Summary

MODS2 Commissioning with some Science Time

Details

UT 04:20 - Starting to clear up, MODS2 closed dome testing underway. Keeping an eye on humidity.

UT 04:36 - Dome open

UT 04:40 - luci.20150513.0001-0005 - 4 sec Dark frames

UT 05:05 - telescope presets don't seem to work - manually movement OK, but otherwise alt and az drives crash.

UT 05:56 - Wondering if I should get out and push

UT 06:28 - Hyperdrive might still be offline, but looks like the source of the problem is the NTP daemon was stopped.

UT 07:13 - And we are open again.

UT 07:20 - Collimating

UT 07:34 - Moving to Telluric Std - seeing 1"

UT 07:42 - Telluric spectra std luci.20150513.0010-0011 Airmass 1.422

UT 07:49 - Moving to SDSSJ1221+1137 Galaxy 1 - Rothberg

UT 08:01 - Starting science spectra guider seeing 1" DIMM 1" (luci.20150513.0015-0019)

UT 08:13 - Seeing is 1.2"

UT 08:25 - Shifting to Galaxy 2 in same field - Airmass 2

UT 08:42 - Starting 2nd galaxy observation (had to nudge by -0.25" in X since very faint) (luci.20150513.0025-0029)

UT 08:44 - Airmass 2.274 - guide 1.2" DIMM 0.66"

UT 09:11 - Done with SDSSJ1221+1137

UT 09:12 - Moving to pointing star, then collimating, before going to FF1532+3242 - mid-z ULIRG - Rothberg

UT 09:17 - Moving to Telluric HD 128039

UT 09:22 - MOS Error - not sure if MASK is in FPU

UT 09:39 - Not the same MOS error as before, unit seems to be in unknown state but never reported moving, clicked initialize, seemed happy, went to move filters, not happy. Several initialization attempts, then called D.Thomspson, who is logged in remotely.

UT 10:51 - D.Thompson recovered MOS

UT 11:01 - Taking flats for calibrations 1.5" slit (first set ending at 0050 too bright, redoing with 6sec DIT (0051-0070)

UT 11:14 - Taking arcs for 1.5" slit (71-74)

UT 11:18 - Finished

-- %USERSIG{BarryRothberg - 2015-05-13}%

Comments

 
Topic revision: r3 - 13 May 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback